# Nonequilibrium steady states in multi-bath quantum collision models

### v1

@@ III Non-Markovian Quantum Collision Models
When using collision models to simulate Markovian dynamics, we make three assumptions regarding the constituents of the bath: (i) they do not interact with each other; (ii) they are initially uncorrelated; (iii) each constituent interacts with $S$ only once. Relaxing these assumptions allows us to extend collision models to simulate non-Markovian dynamics McCloskey and Paternostro (2014); Çakmak et al. (2017); Pezzutto et al. (2016); Zhang et al. (2019); Pleasance et al. (2025).
Here, we introduce non-Markovianity by breaking assumption (i) and introducing intra-bath collisions between consecutive units $A_{n}$ and $A_{n+1}$. We assume that this interaction occurs after the system has collided with $A_{n}$ but before it comes into contact with $A_{n+1}$. In all instances we fix this interaction to be
@@
where $\mathcal{W}$ is as given in Eq. (8), and $0\!\leq\!\delta\!<\!\frac{\pi}{2}$. For the two settings considered previously, the unitary implementing the map is then
@@
The impact of intra-bath interactions is to correlate $S$ with $A_{n+1}$ before they have directly interacted with each other. This leads to a non-Markovian dynamics Breuer et al. (2016) with the parameter $\delta$ controlling the strength of the non-Markovianity McCloskey and Paternostro (2014). For $\delta\!=\!0$ the dynamics reverts to the Markovian case as discussed in Sec. II.3, while $\delta\!=\!\frac{\pi}{2}$ leads to an infinitely non-Markovian dynamics McCloskey and Paternostro (2014) since, for this interaction strength, Eq. (14) becomes a full SWAP and therefore the system is effectively interacting with the same qubit in the environment at every timestep and there is no loss of information.
A useful way to understand the introduction of the intra-bath interactions is through the notion of a memory-depth Campbell et al. (2018). For the Markovian collision models in Sec. II.3, each environmental constituent is involved in only a single collision and, thus, can be traced out after this interaction has taken place since it plays no role in the ensuing dynamics. Clearly, this is not the case for the non-Markovian setting as each $A_{n}$ is involved in two collisions, one with the system and a second with the incoming bath qubit, $A_{n+1}$. Introducing an additional full SWAP operation applied to $A_{n}$ and $A_{n+1}$ in Eqs. (15) allows to recover an effective Markovian dynamics Campbell et al. (2018); Kretschmer et al. (2016a). This is because the additional SWAP operation means that, in effect, the system interacts with the same bath unit, now considered a memory, during every collision. While this may seem like an unnecessary complication, it means that at the level of the system plus memory compound the dynamics is Markovian Campbell et al. (2018); Lorenzo et al. (2017); Kretschmer et al. (2016a). This Markovian embedding of non-Markovian dynamics allows to readily determine the system-memory steady state, from which the steady state of the system directly follows by tracing over the memory’s degrees of freedom.
Our interest is in characterising the steady states arising from the different microscopic models captured by Fig. 1. The ability to introduce non-Markovianity in via intra-environment interactions therefore adds an interesting additional dimension to our study, complementing recent works that focused on the Markovian limit Guarnieri et al. (2020); Prositto et al. (2025). To this end, we aim to assess the impact, if any, that degree of non-Markovianity plays in the resulting steady state properties. Several metrics have been proposed to quantify non-Markovianity Breuer et al. (2016); Addis et al. (2014). We shall focus on the BLP measure Breuer et al. (2009); Laine et al. (2010), which takes an information theoretic approach based on the behaviour of the trace distance between two states,
$D(\rho,\pi)=\frac{1}{2}||\rho-\pi||.$ (16)
@@
where $\Omega_{+}=\bigcup_{i}(a_{i},b_{i})$ is the union of all the time intervals in the dynamics in which the trace distance between $\rho$ and $\pi$ increases, i.e. $\partial_{t}D(\rho(t),\pi(t))>0$. Here, $\rho(t)$ and $\pi(t)$ start from initially distinguishable (orthogonal) states. Equation (17) can be interpreted as summing over all the regions where the trace distance between a pair of system states increases during the dynamics, indicating an increase in their distinguishability and corresponding to an information backflow from the environment. The maximization is taken over all possible pairs of initial states. However, it has been shown that for qubit states, as is the focus of this work, the states that maximise Eq. (17) are always pure, orthogonal states Wißmann et al. (2012). For a collision model which simulates the open system dynamics via discrete time steps, we shall use a discretized version of Eq. (17) such that the non-Markovianity is quantified as McCloskey and Paternostro (2014); Laine et al. (2010),
@@ III.1 Steady State Properties for Non-Markovian Collision Models
We begin by considering how the system approaches the steady state in each setting. For concreteness, we set $\beta\!=\!2$ which fixes initial auxiliary unit states for setting I and the interaction strengths, given in Eq. (9), for setting II. The most relevant parameters dictating the dynamics are then the $S$-$A_{n}$ interaction time $\Delta t$ and the strength of the inter-bath collisions, $\delta$. In Fig. 3 we examine the impact that these parameters have on the ensuing dynamics and the distance from the expected thermal state. We show the fidelity between the evolved state and Eq. (5),
@@
Figure 3: Dynamical approach to the steady state. Upper panel is for setting I and the lower panel is for setting II. We show the fidelity, ($1-F$), against $t\!=\!n\Delta t$, between the evolved system state and Eq. (5) fixing $\beta\!=\!2$. In both settings we consider three sets of parameters: Solid, red lines correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and strong intra-environment interactions $\delta\!=\!0.95\tfrac{\pi}{2}$; Dashed, blue curves correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and weaker intra-environment interactions, $\delta\!=\!0.8\tfrac{\pi}{2}$; Gray, dot-dashed curves correspond to short system-environment collision durations $\Delta t\!=\!0.001$ and strong intra-environment interactions, $\delta\!=\!0.95\tfrac{\pi}{2}$. The insets show the discretized non-Markovianity measure, Eq. (18), evaluated for $\beta\!=\!2$ as a function of intra-environment interactions strength $\delta$. In both main panels the system is initialized in the excited state $\rho_{0}=|1\rangle\langle 1|$. Other parameters used: $\omega=1,\Gamma=1.$
We remark that one could equally employ $\Delta\beta$ as considered in Sec. II.3 since throughout dynamics the system state remains diagonal and hence there exists a $\beta_{e}$. Indeed, we can explicitly relate the fidelity $F$ between $\rho^{\text{th}}$ for a given $\beta$ and $\rho_{n}$ to their temperature difference, $\Delta\beta$,
$F=\frac{\left(1+e^{\tfrac{1}{2}\omega(\Delta\beta+2\beta)}\right)^{2}}{\left(1+e^{\omega\beta}\right)\left(1+e^{\omega(\Delta\beta+\beta)}\right)}$ (20)
@@
Considering first $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ for setting I (blue dashed line) we see that the system monotonically approaches the expected Gibbs state at temperature $\beta=2$. The dynamics appears effectively Markovian, despite the presence of the additional intra-bath interactions. Increasing the intra-bath interactions close to the maximal value is necessary before clear signatures of non-Markovianity emerge, as captured by the red solid curves in Fig. 3 where $\delta\!=0.95\tfrac{\pi}{2}\!$. However, these signatures depend on the value of $\Delta t$. Taking a smaller interaction time we find that, even in the presence of such extremely strong intra-bath interactions, the system’s dynamics still appears Markovian (gray dot-dashed line). Indeed, this can be confirmed by explicitly calculating the BLP measure, Eq. (18), shown in the inset. We see that for a given value of $\Delta t$, there is a critical value of $\delta$ below which the BLP measure is zero, and hence the dynamics appears Markovian. These dynamical differences notwithstanding, for all parameter choices in setting I the system is invariably driven to the expected thermal state. This can be readily confirmed by performing the Markovian embedding and calculating the steady state of system and memory Campbell et al. (2018). The nature of the system-bath interactions necessarily leads to homogenisation and, as will be relevant in later discussions, an asymptotic product state between system and bath units Cusumano et al. (2018).
@@
Turning our attention to setting II, for the same parameters we find some notably different behaviors shown in the lower panel of Fig. 3. For $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ (blue dashed line), the approach to the steady state is still monotonic, while for $\delta\!=\!0.95\tfrac{\pi}{2}$ (red line) the fidelity shows sizeable oscillations before settling to the steady state value. Thus, inline with setting I, strong intra-bath interactions are required to clearly signal the non-Markovian nature of the dynamics. We again find that there is a minimum value of $\delta$ below which the non-Markovian nature of the dynamics is not captured by revivals in the trace distance. We remark that while the BLP measure indicates that setting II is “more” non-Markovian, in the sense that smaller intra-bath collision strengths are necessary for revivals to be present, the overall trend is similar for the two settings and the amount of non-Markovianity as quantified by Eq. (18) is comparable. More interestingly, in the presence of such strong intra-bath interactions we see that the system’s steady state is substantially different from the expected thermal state. From the Sec. II.3 we know that even in the Markovian limit, setting II only thermalises to temperature $\beta$ precisely in the limit of $\Delta t\!\to\!0$, although the deviation is minor for sufficiently small, but strictly finite, values of $\Delta t$ (see Fig. 2). Here, however, we see that the non-Markovian dynamics leads to a significantly larger deviation from the expected thermal state.
@@
We can conclude that settings I and II are essentially equivalent, in the sense that they yield the same thermal steady state, only in the Markovian case $\delta\!=\!0$, and in the limit of $\Delta t\!\to\!0$. Setting I, by construction, will drive the system to the Gibbs state at inverse temperature $\beta$ for any value of $\Delta t>0$ and $0\!\leq\!\delta\!<\!\tfrac{\pi}{2}$ Cusumano et al. (2018). By contrast, setting II drives the system to a NESS which is in general at a different temperature. The deviation of the NESS in setting II from the canonical thermal state becomes more significant as the degree of non-Markovianity is increased.
@@
Figure 4: Panels (a) and (b): Difference in system temperature reached at the steady state, $\Delta\beta\!=\!{\beta_{e}-\beta}$ considering the impact of increasing the A-A interactions $\delta$ and increasing the S-A collision duration $\Delta t$. Panels (c) and (d): Associated steady state heat flux, $\Delta Q/\Delta t$, into $A_{n}^{(0)}$. Other parameters used: $\omega\!=\!1,\Gamma\!=\!1$.
@@
In Fig. 4 we consider the impact of varying $\delta$ and $\Delta t$ on the resulting steady state in setting II, where we consider two values of temperature, $\beta\!=\!0.5$ and $2$. In panels (a) and (b) we show the difference $\Delta\beta$ between the effective temperature $\beta_{e}$ that the steady state reaches, and the canonical temperature $\beta$ appearing implicitly in Eq. (5), i.e. the Gibbs state that the system reaches in the limit of $\Delta t\!\to\!0$. We see a non-trivial interplay between the impact that the two parameters have on the resulting steady-state. For example, stronger non-Markovianity $\delta$ is required to see significant deviations from the expected thermal state when the collision time $\Delta t$ is small. We see that the regions of largest deviation correspond to values of $\Delta\beta\!<\!0$, i.e. the steady state is at a higher temperature than expected. This is markedly different from the Markovian case shown in Fig. 2 where $\Delta\beta\!>\!0$. Considering a larger value for $\beta$ itself, panel (b), we see a shrinking of the regions of most substantial deviation. These areas are generally pushed to higher values of $\delta$ and $\Delta t$, indicating that the non-equilibrium character of the steady state vanishes in the zero temperature limit. This is consistent with the fact that for $\beta\!\to\!\infty$, setting II reduces to the single bath case where homogenization is guaranteed. This variation in the effective temperature that the steady state reaches, corresponding to a transition from effective heating to cooling, implies that in the non-Markovian two bath setting the system can still reach the expected thermal state, shown by the white dashed lines in Fig. 4(a) and (b) indicating the parameter values where $\Delta\beta\!=\!0$. This demonstrates that while the non-Markovianity of the dynamics is a contributing factor for the system reaching a non-equilibrium steady state, there is no simple relationship between the degree of non-Markovianity and the effective temperature to which the system settles.
In the Markovian case $\delta=0$, Eq. (12) shows that in the limit $\Delta t\to 0$ the steady state of setting II is the canonical Gibbs state with $g_{e}\to g$ and hence $\beta_{e}\to\beta$. Interestingly, in the non-Markovian case, a different effective temperature $\beta_{e}\neq\beta$ can be reached even in the limit $\Delta t\to 0$. As can be appreciated visually from Fig. 4(a) and (b), the blue regions of large temperature deviation are pushed to larger and larger $\delta$ as $\Delta t$ becomes smaller. We therefore have a subtle order-of-limits issue, which can be analyzed by considering the effective temperature of the NESS as we send both $\Delta t\to 0$ and $\delta\to 1$ while keeping their ratio $r=\Delta t/(1-\delta)$ constant and finite. By recasting the problem in terms of variables $r$ and $\Delta t$ we find numerically that substantial deviations $\Delta\beta$ survive the $\Delta t\to 0$ limit for fixed values of $r>1$. However, we were not able to extract the non-Markovian analogue of Eq. (12) analytically. Our analysis shows that non-Markovian interactions have a significant effect on the dynamics in setting II, leading to a NESS with with a temperature different from the canonical temperature, even as $\Delta t\to 0$.
@@
The energy-preserving nature of the interactions means that essentially the same heat current is observed due to all the collisions, i.e. the heat flux into the cooler bath is equal and opposite to the heat flux out of the excited state bath and, furthermore, the same magnitude of flux into and out of the auxiliary units is observed due to the intra-bath collisions. As a representative example, we show the heat flux into $A_{n}^{(0)}$ due to the system-bath interaction in Fig. 4(c) and (d). We see that the non-Markovian case exhibits large heat currents to maintain the NESS. As we will see in the following section, these currents allow for the establishment of strong correlations that elucidate why the system settles to a given effective temperature.

### v2

@@ III Non-Markovian Quantum Collision Models
When using collision models to simulate Markovian dynamics, we make three assumptions regarding the constituents of the bath: (i) they do not interact with each other; (ii) they are initially uncorrelated; (iii) each constituent interacts with $S$ only once. Relaxing these assumptions allows us to extend collision models to simulate non-Markovian dynamics ruariMauro; csenyacsa2022entropy; BarisPRA2017; Pezzutto_Paternostro_2016; Zhang2019; Pleasance_2025.
Here, we introduce non-Markovianity by breaking assumption (i) and introducing intra-environment collisions between consecutive units $A_{n}$ and $A_{n+1}$. We assume that this interaction occurs after the system has collided with $A_{n}$ but before it comes into contact with $A_{n+1}$. In all instances we fix this interaction to be
@@
where $\mathcal{W}$ is the SWAP operator,
$\mathcal{W}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix}$, (11)
and $0\!\leq\!\delta\!<\!\frac{\pi}{2}$. For the two settings considered previously, the unitary implementing the map is then
$U_{n}=\begin{cases}U_{AA}(\delta)U^{\text{I}}_{n},&\text{Single~bath~setting~I}\\ U_{AA}^{(0)}(\delta)U_{AA}^{(1)}(\delta)U^{\text{II}}_{n},&\text{Multiple~bath~setting~II}.\end{cases}$ (12)
The impact of intra-environment interactions is to correlate $S$ with $A_{n+1}$ before they have directly interacted with each other. This leads to a non-Markovian dynamics NmReview with the parameter $\delta$ controlling the strength of the non-Markovianity ruariMauro. For $\delta\!=\!0$ the dynamics reverts to the Markovian case as discussed in Sec. II.3, while $\delta\!=\!\frac{\pi}{2}$ leads to an infinitely non-Markovian dynamics ruariMauro since, for this interaction strength, Eq. (10) becomes a full SWAP and therefore the system is effectively interacting with the same qubit in the environment at every timestep and there is no loss of information.
A useful way to understand the introduction of the intra-environment interactions is through the notion of a memory-depth Bassano. For the Markovian collision models in Sec. II.3, each environmental constituent is involved in only a single collision and, thus, can be traced out after this interaction has taken place since it plays no role in the ensuing dynamics. Clearly, this is not the case for the non-Markovian setting as each $A_{n}$ is involved in two collisions, one with the system and a second with the incoming bath qubit, $A_{n+1}$. Introducing an additional full SWAP operation applied to $A_{n}$ and $A_{n+1}$ in Eqs. (12) allows to recover an effective Markovian dynamics Bassano; AA_to_Composite. This is because the additional SWAP operation means that, in effect, the system interacts with the same bath unit, now considered a memory, during every collision. While this may seem like an unnecessary complication, it means that at the level of the system plus memory compound the dynamics is Markovian Bassano; Composite_CM; AA_to_Composite. This Markovian embedding of non-Markovian dynamics allows to readily determine the system-memory steady state, from which the steady state of the system directly follows by tracing over the memory’s degrees of freedom.
Our interest is in characterising the steady states arising from the different microscopic models captured by Fig. 1. The ability to introduce non-Markovianity via intra-environment interactions therefore adds an interesting additional dimension to our study, complementing recent works that focused on the Markovian limit Guarnieri2020PLA; Segal2025QST. To this end, we aim to assess the impact, if any, that degree of non-Markovianity plays in the resulting steady state properties. Several metrics have been proposed to quantify non-Markovianity NmReview; NMmeasures. We shall focus on the BLP measure BLPmeasure; BLPmeasure2, which takes an information theoretic approach based on the behaviour of the trace distance between two states,
@@
where $\Omega_{+}=\bigcup_{i}(a_{i},b_{i})$ is the union of all the time intervals in the dynamics in which the trace distance between $\rho$ and $\pi$ increases, i.e. $\partial_{t}D(\rho(t),\pi(t))>0$. Here, $\rho(t)$ and $\pi(t)$ start from initially distinguishable (orthogonal) states. Equation (14) can be interpreted as summing over all the regions where the trace distance between a pair of system states increases during the dynamics, indicating an increase in their distinguishability and corresponding to an information backflow from the environment. The maximization is taken over all possible pairs of initial states. However, it has been shown that for qubit states, as is the focus of this work, the states that maximise Eq. (14) are always pure, orthogonal states OptimalNM_states. For a collision model which simulates the open system dynamics via discrete time steps, we shall use a discretized version of Eq. (14) such that the non-Markovianity is quantified as ruariMauro; BLPmeasure2,
@@ III.1 Steady State Properties for Non-Markovian Collision Models
We begin by considering how the system approaches the steady state in each setting. For concreteness, we set $\beta\!=\!2$ which fixes initial auxiliary unit states for setting I and the interaction strengths, given in Eq. (7), for setting II. The most relevant parameters dictating the dynamics are then the $S$-$A_{n}$ interaction time $\Delta t$ and the strength of the intra-environment collisions, $\delta$. In Fig. 3 we examine the impact that these parameters have on the ensuing dynamics and the distance from the expected thermal state. We show the fidelity between the evolved state and Eq. (5),
@@
Figure 3: Dynamical approach to the steady state. Upper panel is for setting I and the lower panel is for setting II. We show the fidelity, (1F1-F), against $t\!=\!n\Delta t$, between the evolved system state and Eq. (5) fixing $\beta\!=\!2$. In both settings we consider three sets of parameters: Solid, red lines correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and strong intra-environment interactions $\delta\!=\!0.95\tfrac{\pi}{2}$; Dashed, blue curves correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and weaker intra-environment interactions, $\delta\!=\!0.8\tfrac{\pi}{2}$; Gray, dot-dashed curves correspond to short system-environment collision durations $\Delta t\!=\!0.001$ and strong intra-environment interactions, $\delta\!=\!0.95\tfrac{\pi}{2}$. The insets show the discretized non-Markovianity measure, Eq. (15), evaluated for $\beta\!=\!2$ as a function of intra-environment interactions strength $\delta$. In both main panels the system is initialized in the excited state $\rho_{0}=|1\rangle\langle 1|$. Other parameters used: $\omega=1,\Gamma=4.$
@@
Considering first $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ for setting I (blue dashed line) we see that the system monotonically approaches the expected Gibbs state at temperature $\beta=2$. The dynamics appears effectively Markovian, despite the presence of the additional intra-environment interactions. Increasing the intra-environment interactions close to the maximal value is necessary before clear signatures of non-Markovianity emerge, as captured by the red solid curves in Fig. 3 where $\delta\!=0.95\tfrac{\pi}{2}\!$. However, these signatures depend on the value of $\Delta t$. Taking a smaller interaction time we find that, even in the presence of such extremely strong intra-environment interactions, the system’s dynamics still appears Markovian (gray dot-dashed line). Indeed, this can be confirmed by explicitly calculating the BLP measure, Eq. (15), shown in the inset. We see that for a given value of $\Delta t$, there is a critical value of $\delta$ below which the BLP measure is zero, and hence the dynamics appears Markovian. These dynamical differences notwithstanding, for all parameter choices in setting I the system is invariably driven to the expected thermal state. This can be readily confirmed by performing the Markovian embedding and calculating the steady state of system and memory Bassano. The nature of the system-bath interactions ultimately leads to homogenisation and, as will be relevant in later discussions, an asymptotic product state between system and bath units GiovannettiPRA2018.
@@
Turning our attention to setting II, for the same parameters we find some notably different behaviors shown in the lower panel of Fig. 3. For $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ (blue dashed line), the approach to the steady state is still monotonic, while for $\delta\!=\!0.95\tfrac{\pi}{2}$ (red line) the fidelity shows sizeable oscillations before settling to the steady state value. Thus, inline with setting I, strong intra-environment interactions are required to clearly signal the non-Markovian nature of the dynamics. We again find that there is a minimum value of $\delta$ below which the non-Markovian nature of the dynamics is not captured by revivals in the trace distance. We remark that while the BLP measure indicates that setting II is “more” non-Markovian, in the sense that smaller intra-environment collision strengths are necessary for revivals to be present, the overall trend is similar for the two settings and the amount of non-Markovianity as quantified by Eq. (15) is comparable. More interestingly, however, we now see that in the presence of such strong intra-environment interactions the system’s steady state is substantially different from the expected thermal state, as captured by the solid red line in the lower panel of Fig. 3 settling to a non-zero value.
@@
It follows then that for sufficiently weak intra-environment interactions, aligning with parameter ranges where suitable measures of non-Markovianity fail to reveal any strong memory effects, settings I and II appear essentially equivalent at the level of the steady states insofar as that they yield the same thermal steady state. Setting I, by construction, will drive the system to the Gibbs state at inverse temperature $\beta$ for any value of $\Delta t>0$ and $0\!\leq\!\delta\!<\!\tfrac{\pi}{2}$ GiovannettiPRA2018. By contrast, in the presence of strong intra-environment interactions setting II drives the system to a NESS which is in general at a different temperature. The deviation of the NESS in setting II from the canonical thermal state becomes more significant as the degree of non-Markovianity is increased.
@@
Figure 4: Panels (a) and (b): Difference in system temperature reached at the steady state, $\Delta\beta\!=\!{\beta_{e}-\beta}$ considering the impact of increasing the A-A interactions $\delta$ and increasing the S-A collision duration $\Delta t$. Panels (c) and (d): Associated steady state heat flux, $\Delta Q/\Delta t$, into $A_{n}^{(0)}$. Other parameters used: $\omega\!=\!1,\Gamma\!=\!4$.
In Fig. 4 we consider the impact of varying $\delta$ and $\Delta t$ on the resulting steady state in setting II, where we consider two values of temperature, $\beta\!=\!0.5$ and $2$. As the steady state is diagonal, we can associate to it an effective temperature, $\beta_{e}$. This is formally defined as follows: as the state is diagonal, it can always be written in the form (5), with an effective parameter $g_{e}$. The effective temperature is then defined as
$\beta_{e}=\frac{1}{\omega}\log\left(\frac{1+g_{e}}{1-g_{e}}\right).$ (17)
@@
In panels (a) and (b) we show $\Delta\beta=\beta_{e}-\beta$, i.e. the difference between the effective temperature $\beta_{e}$ that the steady state reaches and the canonical temperature $\beta$ appearing implicitly in Eq. (5). We see a non-trivial interplay between the impact that the two parameters have on the resulting steady-state. For example, stronger non-Markovianity $\delta$ is required to see significant deviations from the expected thermal state when the collision time $\Delta t$ is small. We see that the regions of largest deviation correspond to values of $\Delta\beta\!<\!0$, i.e. the steady state is at a higher temperature than expected. Considering a larger value for $\beta$ itself, panel (b), we see a shrinking of the regions of most substantial deviation. These areas are generally pushed to higher values of $\delta$ and $\Delta t$, indicating that the non-equilibrium character of the steady state vanishes in the zero temperature limit. This is consistent with the fact that for $\beta\!\to\!\infty$, setting II reduces to the single bath case.
@@
The energy-preserving nature of the interactions means that essentially the same heat current is observed due to all the collisions, i.e. the heat flux into the cooler bath is equal and opposite to the heat flux out of the excited state bath and, furthermore, the same magnitude of flux into and out of the auxiliary units is observed due to the intra-environment collisions. As a representative example, we show the heat flux into $A_{n}^{(0)}$ due to the system-bath interaction in Fig. 4(c) and (d). As we will see in the following section, these currents aide in the establishment of strong correlations that contribute to explaining why the system settles to a given effective temperature.